# 3D Uncertain Implicit Surface Mapping using GMM and GP

### v1

@@ Abstract
In this study, we address the challenge of constructing continuous three-dimensional (3D) models that accurately represent uncertain surfaces, derived from noisy and incomplete LiDAR scanning data. Building upon our prior work, which utilized the Gaussian Process (GP) and Gaussian Mixture Model (GMM) for structured building models, we introduce a more generalized approach tailored for complex surfaces in urban scenes, where four-dimensional (4D) GMM Regression and GP with derivative observations are applied. A Hierarchical GMM (HGMM) is employed to optimize the number of GMM components and speed up the GMM training. With the prior map obtained from HGMM, GP inference is followed for the refinement of the final map. Our approach models the implicit surface of the geo-object and enables the inference of the regions that are not completely covered by measurements. The integration of GMM and GP yields well-calibrated uncertainty estimates alongside the surface model, enhancing both accuracy and reliability. The proposed method is evaluated on the real data collected by a mobile mapping system. Compared to the performance in mapping accuracy and uncertainty quantification of other methods such as Gaussian Process Implicit Surface map (GPIS) and log-Gaussian Process Implicit Surface map (Log-GPIS), the proposed method achieves lower RMSEs, higher log-likelihood values and fewer computational costs for the evaluated datasets.
@@
Mapping, Probability and statistical methods, Laser-based, Uncertainty representation.
@@ I Introduction
ACCURACY and reliability (uncertainty quantification) are both critical for mapping in unknown unstructured urban environments, as they guarantee the uncertainty-aware localization, secure navigation and optimal path-planning of autonomous systems, especially with noisy and incomplete measurements. There are many map representations proposed to obtain an accurate environment model while the uncertainty of their mapping outputs is often neglected. Since large errors can also appear in a generally accurate model due to data errors and incompleteness, it is necessary to quantify the uncertainty of the mapping results as well.
@@
The characteristics of the environment representation greatly affect the downstream uncertainty-aware tasks. Traditional occupancy maps [1] represent the space as discrete structured cells and the occupancy of each cell is modelled independently by ray-casting techniques. While they are widely used in localization and collision avoidance, the inconsistency of independent structures poses the challenge for the incomplete, noisy and sparse LiDAR measurements in urban areas. As an alternative, the continuous implicit model [2, 3, 4] shows its advantages in solving this problem. For example, signed-distance fields (SDF) have been suggested for the accuracy and robustness of surface modelling [5]. SDF also naturally provides the informative distance feature; if the distance field is given, the value can be applied directly in ICP localization instead of expensively searching for closet neighbours [4].
@@
In this work, we aim to build continuous 3D surface models of outdoor urban areas using implicit signed-distance functions, together with uncertainty measures. Probabilistic inferences offer the possibility to continuously model environments with well-calibrated uncertainties. Therefore, Gaussian Mixture Model (GMM) and Gaussian Process (GP) inferences are employed and combined in our approach to estimate the distance field models based on LiDAR point clouds. Additionally, we focus on the uncertainty of mapping outputs and investigate if a proper uncertainty measure can provide a good insight into the mapping process. That means, the map quality is correctly quantified and the poor map estimates can be avoided according to the downstream requirements. The existing methods of continuous mapping currently do not sufficiently address this uncertainty issue.
@@
To generate continuous or dense maps, GP-based mapping [6] has been proposed to pose the spatial correlation to the neighbouring points in 2D/3D space. In the work of [6, 7, 8], GP occupancy maps are explored, where the occupancy states are predicted by GP techniques and ”squashed” into occupancy probability in the range of [0, 1] by a logistic regression, which is more robust with sparse data. To reduce the complexity of computational time, approximation skills like sparse kernel[9], data partition[7] and Bayesian committee machine (BCM)[10] are also used in large-scale scenes. Alternatively, Hilbert maps [11, 12] and Bayesian kernel-based inference [13, 14] are also proposed to search for a comparable mapping accuracy as GP, but with faster prediction time. However, the quality of the approximation is degraded as the price of faster speed [14]. Also, while Hilbert maps can be efficient in making predictions, they typically do not provide as rich information on the uncertainty of those predictions compared to GPs.
@@
With a different representation - SDF, the recent work [4] proposed a GP implicit surface (GPIS) map, where the distance fields are regressed instead of occupancy values. In this work, 2D GP is applied first to obtain the derivative information and 3D GP with derivative observations is followed to infer the surface distances. GPIS accurately estimates the distance fields which are very close to the surface but it loses the accuracy at the space far away from the surface. Log-Gaussian Process implicit surface (Log-GPIS) [15, 16] is then proposed to address this issue by applying the logarithmic transformation to a GP formulation, where the regularised Eikonal equation is satisfied to recover the Euclidean distance field (EDF). To reduce the time and space complexity of GPIS, online distance field priors extracted from geometry features are introduced in a recent work [17]. Instead of a feature extraction approach, our work employed GMMs to capture the priors, which compactly parameterize the point clouds and the uncertainties of measurements are embedded in the covariance matrices of the Gaussian components.
@@
While the GP inference, with its data-driven nature, carves local surface features, GMM favours the generalization of the whole measurement distribution. GMMs have been leveraged to generate a continuous parametric representation for point clouds[18, 19], infer high-fidelity occupancy[20, 21, 22], and point cloud registration[23, 24] in the prior work. Nevertheless, model selection remains a challenge in GMM mapping for 3D raw point clouds, i.e., it is difficult to determine the proper number of components for various complex scenes.
@@
Recently, Neural Radiance Field (NeRF) representation [25] has been proposed to model the environment with a continuous implicit function of radiance and occupancy information, which shows its great potential in mapping, localization, and planning. However, it lacks the sufficient well-calibrated uncertainty representation of the mapping results. Also, due to the fixed architecture of the underlying network, it is not able to directly adapt to the representation requirement of the scene [19].
@@
In our previous work[26], we proposed GP-mapping for building models, with 1D GMM as prior for the surface depth. In each local facade frame, the surface is represented in 2.5D. In this letter, we generalize the previous methods to more complex structures, employing 4D GMMs to obtain the model prior and convert the constrained local 2.5D surfaces to fully 3D representation. Also, the mapping process is simplified since the segmentation of local surfaces is not required anymore. When GMMs are extended to higher dimensions, the issue of model selection (choice of the GMM component number) arises. Therefore, a hierarchical GMM technique is employed to solve the selection of component numbers and speed up GMM training. GP inference is followed to correct irregular parts of GMM priors that do not comply with the original training points.
@@
  • We map uncertain 3D continuous signed-distance fields in urban areas with probabilistic inferences using laser scanning point clouds, which can be applied in various downstream tasks and can be updated probabilistically. The experimental results demonstrate the value of the uncertainty measures, showing that estimations with lower uncertainty lead to superior surface reconstruction outcomes.
  • We expand the existing work[26] to a more generalized approach by applying a higher dimensional hierarchical GMM, combined with GP inferences to yield a fully 3D map with efficient surface prediction and to maintain a faithful uncertainty description. This approach establishes a seamless connection between the HGMM surface distance estimation and the GP framework, contributing to a cohesive and synergistic modeling strategy.
  • We evaluate the accuracy and the faithfulness of the uncertainty quantification with both real-world LiDAR point clouds, collected by the mobile mapping system (MMS) and simulated point clouds from the Carla simulator [27]. The Root mean squared error (RMSE) and averaged log-likelihood are used as metrics to evaluate the results of GPIS, Log-GPIS, and our method (GMM alone and GMMGP). The potential of uncertainty measures is also discussed.
The rest of the letter is structured as follows. Section II introduces the map representation being studied. Section III details the HGMM and its regression techniques. Following this, Section IV explains GP inference, incorporating the HGMM as a prior. Section V presents the experimental evaluation of our approach, utilizing real-world LiDAR data, and offers both qualitative insights and a quantitative comparison against state-of-the-art benchmarks. Conclusions and an outlook for future research are provided in Section VI. Figure 1 illustrates the overall workflow of the mapping process.
@@ II Map representation
It should be noted that in scenarios where the sign information is disregarded, one obtains the Euclidean Distance Field (EDF) as a result. EDF defines spatial relationships purely based on the magnitude of distances, absent the directional or inside-outside context provided by signed distance measures. The surface representation can be regressed by both the GMM and GP approaches and both provides the probabilistic uncertainty for the distance estimates. The GMM results serve as dynamic prior mean and kernel functions in the GP training.
@@ III-A Hierarchical Gaussian Mixture Models
A standard Gaussian Mixture Model is composed of $K$ Gaussian components, described by a set of parameters $\Theta=\{(\pi_{k},\bm{\mu_{k}},\bm{\Sigma_{k}})\}^{K}_{k=1}$, as illustrated in:
@@
where $\pi_{k}$ is the prior weight, ${\bm{\mu_{k}}}$ and ${\bm{\Sigma_{k}}}$ are the Gaussian parameters. Given a perfect $K$, the components will properly approximate the object surface. Assuming the surface can be decomposed with many planar patches, each component estimates one planar patch, and the 3-dimensional manifold degenerates into a nearly 2-dimensional structure. This implies that performing an eigenvalue decomposition on the covariance matrices of the respective components results in their smallest eigenvalues approaching zero.
However, a 3-dimensional GMM cannot completely degenerate into 2D. Instead, what can be achieved is the optimization of the model by targeting very small values for the principal curvatures or the eigenvalues associated with the smallest dimension. It is non-trivial to select the proper number of GMM components to achieve this goal. The idea of HGMM is to create a hierarchical tree structure for GMM models. This structure initiates modeling with a small number of Gaussian nodes to encapsulate the data. Subsequently, the complexity of each node is evaluated to select the nodes that should be split. Here, their complexity is assessed based on the deviation from a near-2D (or ’flat’) structure. Nodes exhibiting complexity beyond a predefined threshold are then subdivided into more nuanced sub-nodes at the next level of the hierarchy; i.e., new Gaussian components and GMM training are required for them. In this case, the complexity of the Gaussian node is quantified by principal curvature, given by:
@@
where $\lambda_{i}$ is the $i-$th largest eigenvalue of the principal components, $\lambda_{n}$ is the smallest eigenvalue. If $c$ is small, the structure is relatively flat.
@@
(d) Results
Figure 2: HGMM results of an example building data: the error ellipsoids of Gaussian components are visualized. (a) is the original measurements. (b) shows four initial Gaussian components, and they split into the next level in (c). The new components are checked by the principal curvature to see if they should be further split. (d) presents the final GMM models.
@@ III-B Gaussian Mixture Regression
To infer the surface distance with HGMM, the original 3D HGMM is extended to 4D space and the Gaussian mixture regression (GMR) technique is employed. Given LiDAR measurements ${\bf{X}}=\{{\bm{x}}_{i}\in\mathbb{R}^{3}\}_{i=1}^{N}$, the 3D points are appended by a fourth dimension representing surface distances, denoted as $\bm{Y}=\{y_{i}\}_{i=1}^{N}$, forming the complete dataset for training: $\bm{D}=(\bf{X},\bm{Y})$. Following the EM training, the parameters $\Theta=\{(\pi_{k},\bm{\mu_{k}},\bm{\Sigma_{k}})\}^{K}_{k=1}$ are derived and each $k$-th component is a jointly Gaussian, given by:
@@
In GMM regression, the rule of conditionals of multivariate Gaussian distributions is applied to estimate the queried surface distance conditioned on the 3D point coordinates. For each component, the posterior conditional distribution reads:
@@
The overall surface regression is the mixture of different Gaussian components. In fact, not all components in GMM have effective impacts on the estimation of a queried area, rather, only some neighboring Gaussians are important, and we call these components the active components. Thus, $J$ components will be selected as active components according to the marginal probability density $p_{j}(\bm{x_{*}}|\bm{\mu_{jX}},\bm{\Sigma_{\bm{jXX}}})$.
@@
As illustrated above, the signed distance value $y$ of each data point is required in the training process. However, the measurements are all collected when the scanning ray hits the surface, i.e., $y=0$ for all measurements. To obtain valid covariance between spatial coordinates and surface distances, the original dataset is augmented with virtual points along the normal direction of the hit points and at a small distance from the surface. Implement details will be explained in Section V.
Despite the extension of the hierarchical model from 3D to 4D and the consequent changes in the interpretations of complexity and split criteria, these criteria remain applicable and suitable in the 4D context. When the smallest eigenvalue approaches zero, it indicates that the variance along one principal axis is minimal. The N-dimensional manifolds are nearly degenerated into (N-1)-dimensional manifolds. This condition typically leads to a tighter correlation between the variables involved. Consequently, in our case, the regression of the fourth dimension (surface distances) based on point coordinates will likely exhibit less noise and can be expected to be more accurate. In the meantime, it often forces the first 3D variables to have a more flat structure as well.
Although the hierarchical strategy helps us to adaptively find the number of Gaussian components and quickly optimize them, the performance is still sensitive to the thresholds of the complexity and the size of the modeling objects. Using a too-big threshold sometimes ignores the details of large objects, while it might oversplit the nodes with a too-small threshold. Also, one single threshold may affect the ability to fit the scene with both large and small objects and also the threshold can become too small with a higher hierarchy of smaller nodes. Therefore, HGMM is an approximate of the ideal GMM with the best fidelity, which can be further improved. Rather than selecting the perfect number of components, GP-based mapping is followed to refine the results from HGMM. More specifically, parameters optimized in HGMM are used as priors for GP training, which will be introduced in Section IV.

### v4

@@ Abstract
In this study, we address the challenge of constructing continuous three-dimensional (3D) models that accurately represent uncertain surfaces, derived from noisy and incomplete LiDAR scanning data. Building upon our prior work, which utilized the Gaussian Process (GP) and Gaussian Mixture Model (GMM) for structured building models, we introduce a more generalized approach tailored for complex surfaces in urban scenes, where GMM Regression and GP with derivative observations are applied. A Hierarchical GMM (HGMM) is employed to optimize the number of GMM components and speed up the GMM training. With the prior map obtained from HGMM, GP inference is followed for the refinement of the final map. Our approach models the implicit surface of the geo-object and enables the inference of the regions that are not completely covered by measurements. The integration of GMM and GP yields well-calibrated uncertainties alongside the surface model, enhancing both accuracy and reliability. The proposed method is evaluated on real data collected by a mobile mapping system. Compared to the performance in mapping accuracy and uncertainty quantification of other state-of-the-art methods, the proposed method achieves lower RMSEs, higher log-likelihood values and lower computational costs for the evaluated datasets.
@@
Mapping, Laser-based, Probability and statistical methods, Uncertainty representation.
@@ I Introduction
ACCURACY and the possibility to detect gross errors are both critical for mapping in unknown unstructured urban environments, as they guarantee uncertainty-aware localization, secure navigation and optimal path-planning of autonomous systems, especially with noisy and incomplete measurements. Some popular mapping frameworks [1, 2, 3] emphasize accuracy and efficiency while the uncertainty of mapping outputs is not sufficiently addressed. Since large errors can also appear in a generally accurate model due to data errors and incompleteness, it is necessary to quantify the uncertainty of the mapping results as well.
@@
The characteristics of the environment representation greatly affect the downstream uncertainty-aware tasks. Traditional occupancy maps [1] represent the space as discrete structured cells and the occupancy of each cell is modelled independently by ray-casting techniques. While they are widely used in localization and collision avoidance, the inconsistency of independent structures poses the challenge for the incomplete, noisy and sparse LiDAR measurements in urban areas. As an alternative, continuous implicit models [5, 6, 4, 7, 8, 9, 10] show their advantages in solving this problem. For example, signed-distance fields (SDF) have been suggested to improve the accuracy and robustness of surface modelling [7].
@@
In this work, we aim to build continuous 3D surface models of outdoor urban areas using implicit signed-distance functions, together with uncertainty measures. Probabilistic inferences offer the possibility to continuously model environments with well-calibrated uncertainties. Therefore, Gaussian Mixture Model (GMM) and Gaussian Process (GP) inferences are employed and combined in our approach to estimate the distance fields in large-scale scenes, especially with urban manmade structures, based on LiDAR point clouds. Additionally, we focus on the uncertainty of mapping outputs and investigate if a proper uncertainty measure can provide a good insight into the mapping process. That means the map quality is correctly quantified and poor map estimates can be avoided accordingly. The existing methods of continuous mapping currently do not sufficiently address this uncertainty issue.
@@
To generate continuous maps, GP-mapping [11] introduces the spatial correlation to the neighbouring points. In GP occupancy maps [11, 12, 13], the occupancy states are predicted by GPs and ”squashed” into probability in the range of [0, 1] by a logistic regression, which is robust with sparse data. To reduce the computation time, approximation methods like sparse kernel[14] and data partition[12] are used in large-scale scenes. Alternatively, Hilbert maps [15] and Bayesian kernel inference [16] search for a comparable accuracy as GP, but with faster prediction time. However, the quality of the approximation is degraded at the price of faster speed [16].
@@
Lee et al. [4] proposed a GP implicit surface (GPIS) map, where the distance fields are regressed instead of occupancy values. In this work, 2D GP is applied first to obtain the derivatives and 3D GP with derivative observations is followed to infer the surface distances. GPIS accurately estimates the distance fields which are close to the surface but it loses the accuracy at the regions far from the surface. Log-GPIS [8] is then proposed to address this issue by applying the logarithmic transformation to a GP formulation, but it faces the trade-offs between accuracy and interpolation abilities. Le Gentil et al. [9] solved this problem by leveraging the relation between the kernel and the distance fields. However, the training time of those methods still scales with the increasing data size. To reduce the time/memory complexity of GPIS, distance field priors extracted from geometry features are introduced in [10]. Instead of a feature extraction, this paper employed GMMs to capture the priors, which compactly parameterize point clouds, and the uncertainties are embedded in the covariance matrices of the Gaussian components.
@@
While the GP inference, with its data-driven nature, carves local surface features, GMM favours the generalization of the whole measurement distribution. GMMs have been leveraged to generate a continuous parametric representation for point clouds[17, 18, 19, 20, 21], infer high-fidelity occupancy[22, 23, 24, 25], and point cloud registration[26, 27] in the prior work. Nevertheless, model selection remains a challenge in GMM mapping for 3D point clouds[18], i.e., it is difficult to determine the proper number of components for various complex scenes. Information criterion-based model selection is time-intensive as it tests various numbers to find the elbow point. Recent studies have introduced adaptive strategies that employ hierarchical structures [27] and information-theoretic optimization [18]. However, hierarchical methods remain sensitive to hyper-parameters involved in splitting and encounter trade-offs between accuracy and efficiency [27], while SOGMM [18] may face the issue of selecting an optimal bandwidth parameter with the limited compute in complex outdoor scenarios.
@@
A neural network-based method - Neural Radiance Field (NeRF) [2] has been proposed to model the environment with a continuous implicit function of radiance and occupancy, which shows its great potential in mapping, localization, and planning. However, it lacks the sufficient well-calibrated uncertainties of the mapping results. Also, due to the inherent neural components, obtaining high-quality results continues to incur substantial expenses in training and rendering. The fixed architecture of the underlying network makes it unable to directly adapt to different scenes[18].
Recently, Gaussian splatting (GS)[3] leverages the 3D Gaussians as point-based blending with their positions and opacity, efficiently and flexibly rendering the color information compared to NeRF. Unlike GMM-mapping, GS is not explicitly calibrated to represent the point cloud distribution. Gaussians do not precisely match the actual surface [28]. The geometry information is modelled by rendering depth images and requires additional enhancement, e.g. optical flow loss in [29]. Also, the inherent uncertainty of Gaussians in GS is not investigated.
@@
In our previous work[30], we proposed GP-mapping focusing on structured buildings, with 1D GMM as prior for the surface depth. In each local facade frame, the surface is represented in 2.5D. In this paper, we generalize the previous methods to more complex structures, employing 4D HGMMs to obtain the model prior and convert the local 2.5D surfaces to fully 3D representations, with fewer geometrical constraints for more general shapes. Also, the mapping process is simplified since the segmentation of local surfaces is not required anymore, reducing training time and minimizing potential segmentation errors. When GMMs are extended to higher dimensions, the issue of model selection (choice of the component number) arises. Therefore, a hierarchical GMM technique is employed to solve the selection of component numbers and speed up GMM training. GP inference is followed to correct irregular parts of GMM priors that do not comply with the original training points.
@@
  • We map uncertain 3D continuous SDFs in urban areas with probabilistic inferences using laser scanning point clouds, which can be applied in various downstream tasks and can be updated probabilistically. The experimental results demonstrate the value of the uncertainty measures, showing that estimations with lower uncertainty lead to superior surface reconstruction outcomes.
  • We expand the prior work[30] to a more generalized approach by applying a higher dimensional hierarchical GMM, combined with GPs to yield a fully 3D map with efficient surface prediction and to maintain a uncertainty description. This approach establishes a seamless connection between the HGMM surface distance estimation and the GP framework, contributing to enhanced accuracy, efficiency and the robustness to more general shapes.
  • We evaluate the accuracy and the faithfulness of the uncertainty quantification with real-world LiDAR point clouds, collected by the mobile mapping system (MMS). The Root mean squared error (RMSE) and averaged predictive log-likelihood are used as metrics to evaluate the results.
Figure 1 illustrates the overall mapping workflow, showing how HGMM is integrated into GP, with detailed explanations provided in the subsequent sections.
@@ II Map representation
It should be noted that in scenarios where the sign information is disregarded, one obtains the Euclidean Distance Field (EDF) as a result. EDF defines spatial relationships purely based on the magnitude of distances, absent the directional or inside-outside context provided by signed distance measures. The surface representation can be regressed by the GMM and GP approaches and both provide the probabilistic uncertainty for the distance estimates. The GMM results serve as non-stationary prior mean and kernel functions in the GP training.
@@ III-A Hierarchical Gaussian Mixture Models
A standard GMM is composed of $K$ Gaussian components, described by a set of parameters $\Theta=\{(\pi_{k},\bm{\mu_{k}},\bm{\Sigma_{k}})\}^{K}_{k=1}$, as given in:
@@
where $\pi_{k}$ is the prior weight, ${\bm{\mu_{k}}}$ and ${\bm{\Sigma_{k}}}$ are the Gaussian parameters. Given a perfect $K$, the components will properly approximate the object surface. Assuming the surface can be decomposed with many planar patches, each component estimates one planar patch, and the 3-dimensional manifold degenerates into a nearly 2D structure. This implies that performing an eigenvalue decomposition on the covariance matrices of the respective components results in their smallest eigenvalues approaching zero.
However, a 3D GMM cannot completely degenerate into 2D. Instead, what can be achieved is the optimization of the model by targeting very small values for the principal curvatures or the eigenvalues associated with the smallest dimension. It is non-trivial to select the proper number of GMM components to achieve this goal. The idea of HGMM is to create a hierarchical tree structure for GMM models. This structure initiates modelling with a small number of Gaussian nodes to encapsulate the data. Subsequently, the complexity of each node is evaluated to select the nodes that should be split. Here, their complexity is assessed based on the deviation from a near-2D (or ’flat’) structure. Nodes exhibiting a complexity beyond a predefined threshold are then subdivided into more nuanced sub-nodes at the next level of the hierarchy; i.e., new Gaussian components and GMM training are required for them. In this case, the complexity of the Gaussian node is quantified by the principal curvature, given by:
@@
where $\lambda_{i}$ is the $i$-th eigenvalue of the principal components, $\lambda_{n}$ is the smallest eigenvalue. If the principal curvature is small, the structure is relatively flat. The threshold is set to the same empirical value as the previous work [17].
Figure 2 visualizes an example where GMM structures gradually split at hierarchical levels. In the lower right regions of the building, where there are some protruded balconies, Gaussian components split from one green Gaussian (level 1) into more smaller Gaussians (level 3) to fit the local smaller surface structures.
@@
(d) Level 3
Figure 2: Visualization of the error ellipsoids of Gaussian components in HGMM. (a) is the data. (b) shows four initial Gaussians, and they split into the next level in (c). The new components are checked by the principal curvature to see if they should be further split. (d) presents the final models.
@@ III-B Gaussian Mixture Regression
To infer the surface distance with HGMM, the original 3D HGMM is extended to 4D space and the Gaussian mixture regression (GMR) technique is employed. Given LiDAR measurements ${\bf{X}}=\{{\bm{x}}_{i}\in\mathbb{R}^{3}\}_{i=1}^{N}$, the 3D variables (coordinates) are appended by a fourth dimension representing surface distances, denoted as $\bm{Y}=\{y_{i}\}_{i=1}^{N}$, forming the complete dataset for training: $\bm{D}=(\bf{X},\bm{Y})$. Following the EM training, the parameters $\Theta=\{(\pi_{k},\bm{\mu_{k}},\bm{\Sigma_{k}})\}^{K}_{k=1}$ are derived and each $k$-th component is a jointly Gaussian, given by:
@@
In GMM regression, the rule of conditionals of multivariate Gaussian distributions is applied to estimate the queried surface distance conditioned on the 3D coordinates. For each component, the posterior conditional distribution reads as:
@@
The overall surface regression is the mixture of different Gaussian components. In fact, not all components in the GMM have effective impacts on the estimation of a queried area, rather, only some neighboring Gaussians are important, and we call these components the active components. Thus, $J$ components will be selected as active components according to the marginal probability density $p_{j}(\bm{x_{*}}|\bm{\mu_{jX}},\bm{\Sigma_{\bm{jXX}}})$.
@@
As illustrated above, the signed distance value $y$ of each data point is required in the training process. However, the measurements are all collected when the scanning ray hits the surface, i.e., $y=0$ for all measurements. To obtain valid covariances between spatial coordinates and surface distances, the original dataset is augmented with virtual points along the normal direction of the hit points and at a small distance from the surface. The extension of GMM dimension for map regression was originally proposed in the prior work [22]. The key differences are: 1) the proposed method applies a “top-down” strategy with faster computation speed; 2) SDF is regressed instead of occupancy.
Despite the extension of the hierarchical model from 3D to 4D and the consequent changes in the interpretations of complexity and split criteria, these criteria remain applicable and suitable in the 4D context. When the smallest eigenvalue approaches zero, it indicates that the variance along one principal axis is minimal. The N-dimensional manifolds are nearly degenerated into (N-1)-dimensional manifolds. This condition typically leads to a tighter correlation between the variables involved. Consequently, in our case, the regression of the fourth dimension (surface distances) based on point coordinates will likely exhibit less noise and can be expected to be more accurate. In addition, it often forces the first 3D variables to have a more flat structure as well.
Although the hierarchical strategy helps us to adaptively find the number of Gaussian components and quickly optimize them, the performance is still sensitive to the thresholds of the complexity and the size of the modeling objects. Using a too-big threshold sometimes ignores the details of large objects, while it might oversplit the nodes with a too-small threshold. Also, one single threshold may affect the ability to fit the scene with both large and small objects and also the threshold can become too small with a higher hierarchy of smaller nodes. Therefore, HGMM is an approximation of the ideal GMM with the best fidelity, which can be further improved. Rather than selecting the perfect number of components, GP-based mapping is followed to refine the results from HGMM. More specifically, parameters optimized in HGMM are used as priors for GP training, which will be introduced in Section IV.